A toy truck of mass 10 kg


1
Expert's answer
2021-02-28T07:35:14-0500

The acceleration of centre of mass of the two is


"a_{cm}=\frac{m_1a_1+m_2a_2}{m_1+m_2}\\a_{cm}=\frac{20(10)+10(40)}{10+20}=20\frac{m}{s^2}"